OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) If a triangle has an area of 56 sq. inches, with base of 9 inches longer that its altitude, find the length of its base.

OpenStudy (anonymous):

 Area = 1/2 base x height.

OpenStudy (anonymous):

(1/2)(b)(b+9)=56 (b)(b+9)=108 now use the quadratic formula

OpenStudy (anonymous):

\[x=\frac{-9\pm \sqrt{513}}{2}\]

OpenStudy (anonymous):

It's worth remembering that only positive roots will be valid (since lengths cannot be negative)

OpenStudy (anonymous):

@Dalvoron, good point